In a triangle ABC, 2ac sin 1/2 (A – B + C) =
(A) a2+b2-c2                                              (B) c2+a2-b2
(C) b2-c2-a2                                               (D) c2-a2-b2

Grade:12

1 Answers

Ramesh V
14 years ago

A+B+C=Pi

so A+C = Pi - B

so here it will be :  2ac.sin(Pi/2 -B) = 2ac.cosB

by cosine rule, we have: c2+a2-b2 = 2ac cos B

option B

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh 
IIT Kgp - 05 batch